# When a Reinforcement Learning Agent Encounters Unknown Unknowns

### v1

@@ Appendix D When a New Observation Hurts Awareness Confidence – A Reverse Bayesian Perspective
$\frac{1}{|\mathcal{S}_{t+1}-\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t+1}-\mathcal{S}_{t}}\tilde{\overline{V}}_{h}^{t}(\hat{s})-V^{*}_{h}(\hat{s})$ $\geq\frac{1}{|\mathcal{S}_{t+1}-\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t+1}-\mathcal{S}_{t}}[\frac{1}{|\mathcal{S}_{t}|}\sum_{s^{\prime}\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(s^{\prime})-V^{*}_{h}(s^{\prime})]$
@@ Appendix E Proof of Corollary 1
Appendix F Proof of Theorem 1
Proof of Theorem 1.
We prove the lemma under Assumption 1. For $s^{\prime}\in\mathcal{S}_{0}$, we have $t(s^{\prime})=0\leq\sqrt{T}$. For $s^{\prime}\in\mathcal{S}-\mathcal{S}_{0}$, we have $\forall(s,a)\in\mathcal{S}\times\mathcal{A},P(s^{\prime}|s,a)\geq 1-(\frac{\delta}{3})^{\frac{1}{H[\sqrt{T}]}}\triangleq\theta$. Thus, the following holds $\mathbb{P}(s^{\prime}\notin O_{[\sqrt{T}]})\leq(1-\theta)^{H[\sqrt{T}]}$. From the relationship between $O_{t}$ and $\mathcal{S}_{t}$, we get
$\mathbb{P}(s^{\prime}\notin\mathcal{S}_{[\sqrt{T}]})\leq(1-\theta)^{H[\sqrt{T}]}=\frac{\delta}{3}.$
Finally, we get $t(s^{\prime})\leq[\sqrt{T}]\leq\sqrt{T}$ with probability at least $1-\frac{\delta}{3}$. ∎
Appendix G Proof of Lemma 3 and Lemma 4
@@
Appendix H Concentration Events
@@
[…]
$|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h}^{k-1})(s,a)}+6H\zeta\},$
[…]
@@
[…]
@@
[…]
@@
[…]
@@
For $\mathcal{E}^{v_{2}}$, based on Lemma 16, we have $Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)\leq 2H^{2}Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$. Based on $\zeta=log(\frac{32e(2T+1)}{\delta})$ and Theorem 6, we get
@@
$|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}^{k-1}_{h,s,a})(s,a)|\leq\frac{1}{4}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H\zeta,$
@@
$|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)|\leq\frac{1}{4}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+44H^{2}\zeta.$
@@
[…]
@@
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)|\leq\sum_{k=t(s)}^{t}\overline{p}_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)+8H^{2}\zeta\},$
$\mathcal{G}^{v_{1}}\triangleq\{\forall t\in[T],\forall h\in[H],\forall(s,a)\in\mathcal{S}_{t}$$\times\mathcal{A}:|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}})(s,a)|$
@@
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s)(\mathring{\overline{V}}_{h}^{k-1}-V_{h}^{\pi^{t}})(s)|\leq\frac{1}{4H}\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s)|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s)+14H^{2}\zeta\},$
@@
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)|$ $\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\overline{p}_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)^{2}}+6H^{2}\zeta$
@@
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s)(\mathring{\overline{V}}_{h}^{k-1}-V_{h}^{\pi^{t}})(s)|$ $\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s)|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s)^{2}}+6\zeta H^{2}$
@@ Appendix I Optimism
$=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)]-p_{h}\mathring{V}_{h,s,a}^{t}(s,a)$
$=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)]$
@@
$\leq\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}.$
@@
$W_{h}^{t}(s,a)=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}.$
@@
$\leq\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H^{3}\frac{log(T)\zeta}{n_{h}^{t}(s,a)}.$
@@
$\ \ \ \ \ |(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}|$
$\leq|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}+p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|\cdot|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|$
@@
$W_{h}^{t}(s,a)$ $=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
$\geq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
@@
$\geq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(p_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
@@
$\beta_{h}^{t}(s,a)$ $\geq 2\sqrt{(W_{h}^{t}(s,a)+88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)})\frac{\zeta}{n_{h}^{t}(s,a)}}+(53-2\sqrt{88}-14)H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
@@
$\geq 2\sqrt{\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+20H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
@@
$\mathring{\overline{V}}_{h}^{t}(s)$ $=\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(\hat{s})$
@@
Appendix J Proof of Theorem 2

### v2

@@ Appendix D When a New Observation Hurts Awareness Confidence – A Reverse Bayesian Perspective
[…]
$\geq\frac{1}{|\mathcal{S}_{t+1}-\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t+1}-\mathcal{S}_{t}}[\frac{1}{|\mathcal{S}_{t}|}\sum_{s^{\prime}\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(s^{\prime})-V^{*}_{h}(s^{\prime})]$
@@ Appendix E Proof of Corollary 1
Appendix F Proof of Lemma 2 and Lemma 3
@@
Appendix G Concentration Events
@@
$\mathcal{E}^{v_{1}}\triangleq\{$ $\forall t\in[T],\forall h\in[H],\forall(s,a)\in\mathcal{S}_{t}\times\mathcal{A}:$
$|\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|$
$\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h}^{k-1})(s,a)}+6H\zeta\},$
$\mathcal{E}^{v_{2}}\triangleq\{$ $\forall t\in[T],\forall h\in[H],\forall(s,a)\in\mathcal{S}_{t}\times\mathcal{A}:$
@@
$\mathcal{E}^{m_{\omega}}\triangleq\{$ $\forall t\in[T],\forall h\in[H],\forall(s,a)\in\mathcal{S}_{t}\times\mathcal{A}:$
@@
$\mathcal{E}^{m}\triangleq\{$ $\forall t\in[T],\forall h\in[H],\forall(s,a)\in\mathcal{S}_{t}\times\mathcal{A}:$
@@
$\forall s,s^{\prime}\in\mathcal{S},\ \forall t\in[T],\ 0\leq\mathring{\overline{V}}_{h+1}^{t}(s)\leq H,\ 0\leq\mathring{V}^{t}_{h,s,a}(s^{\prime})-\mathring{\overline{V}}_{h+1}^{t}(s^{\prime})\leq H.$
@@
For $\mathcal{E}^{v_{2}}$, based on Lemma 16, we have
$Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)\leq 2H^{2}Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a).$
Based on $\zeta=log(\frac{32e(2T+1)}{\delta})$ and Theorem 5, we get
@@
$\leq\frac{1}{4}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)+14H\zeta,$
@@
$\mathring{\gamma}_{h}^{t}(s,a)\leq H,\ 0\leq\mathring{V}_{h,s,a}^{t-1}(s^{\prime})-\mathring{V}_{h+1}^{t-1}(s^{\prime})\leq H,\ \sqrt{xy}\leq x+y,$
@@
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)|$
$\leq\sum_{k=t(s)}^{t}\overline{p}_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)+8H^{2}\zeta\},$
$\mathcal{G}^{v_{1}}\triangleq\{$ $\forall t\in[T],\forall h\in[H],\forall(s,a)\in\mathcal{S}_{t}\times\mathcal{A}:$
@@
$|\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s)(\mathring{\overline{V}}_{h}^{k-1}-V_{h}^{\pi^{t}})(s)|$
$\leq\frac{1}{4H}\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s)|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s)+14H^{2}\zeta\},$
@@
$\ \ \ \ |\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)|$
$\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\overline{p}_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t}})(s,a)^{2}}+6H^{2}\zeta$
@@
$\ \ \ \ |\sum_{k=t(s)}^{t}(\chi_{h}^{k}-\overline{p}_{h}^{k})(s)(\mathring{\overline{V}}_{h}^{k-1}-V_{h}^{\pi^{t}})(s)|$
$\leq\sqrt{2\zeta\sum_{k=t(s)}^{t}\overline{p}_{n}^{k}(s)|\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{t}}|(s)^{2}}+6\zeta H^{2}$
@@ Appendix H Optimism
$=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)]$
$\ \ \ \ -p_{h}\mathring{V}_{h,s,a}^{t}(s,a)$
$=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)$
$\ \ \ \ +\mathring{\gamma}_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)]$
@@
$\leq\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
$\ \ \ \ +14H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}.$
@@
$W_{h}^{t}(s,a)=$ $\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)$
$-\left(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)\right)^{2}.$
@@
$\leq\frac{1}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}^{k-1}_{h,s,a}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
[…]
@@
$\ \ \ \ |(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$ (5)
$\ \ \ \ -(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}|$
$\leq|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}+p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|$ (6)
$\ \ \ \ \cdot|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(\mathring{p}_{h}^{k}-p_{h})\mathring{\overline{V}}_{h+1}^{k-1}(s,a)|$
@@
$W_{h}^{t}(s,a)$ $=\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)$ (8)
$\ \ \ \ -(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
$\geq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)$ (9)
$\ \ \ \ -(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
@@
$\geq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)$ (10)
$\ \ \ \ -\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)(p_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
@@
$\beta_{h}^{t}(s,a)$ $\geq 2\sqrt{(W_{h}^{t}(s,a)+88H^{2}\frac{\zeta}{n_{h}^{t}(s,a)})\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\ \ \ \ +(53-2\sqrt{88}-14)H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
@@
$\geq 2\sqrt{\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(\mathring{\overline{V}}_{h+1}^{k-1})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
$\ \ \ \ +20H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
@@
[…]
@@
Appendix I Proof of Theorem 1